OpenStudy (anonymous):

Find an equation of a line tangent to a circle x^2+y^2=10x^2+y^2=10 at (3,1). A) y=1/3x B) y = -3x + 10 C) y=1/3x+10/3 D) y = -3x + 1

OpenStudy (ash2326):

Is this the equation of circle \[x^2+y^2=10x^2+y^2=10\] ?

OpenStudy (anonymous):

Equation of circle is x^2 + y^2 = 10. So, the center is (0,0). Now, you have to find the equation of the line passing between (0,0) and (3,1). Then, find a line perpendicular to this line passing through (3,1).

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Those are the steps. Tangent = Line perpendicular to the line connecting to the center.

OpenStudy (anonymous):

I dont know any of this. im home schooled lol

OpenStudy (anonymous):

my parents are highschool drop outs, im on my own

OpenStudy (anonymous):

You just need to take it one step at a time and breakdown this problem into three problems: 1) Find center of circle. Easy part given the equation it is (0,0). 2) Find equation of a line passing through two points (you should be able to find the procedure anywhere online) 3) Find equation of line perpendicular to the line in step #2 (again, steps and procedures available online)

OpenStudy (anonymous):

I think its B. Is that right?

OpenStudy (anonymous):



OpenStudy (anonymous):

This diagram should help clarify the tangent and equations needed.

OpenStudy (anonymous):

B is correct. But, again, more important how you get there.